Ang company that produces pleasure boats has decided to expand one of its lines. The company is considering three Location alternatives. The cost structure for each location is given as shown in the table.

Location

Fixed Cost/yr

variable cost/unit

L1

7,958

3.1

L2

5,879

4.5

L3

11,897

2.2


a) The range of output for location alternative 1 that would yield the lowest total cost is ______   (Ex. In whole number, use Q, <, <=, >, >= to specify the range)

b) The range of output for location alternative 2 that would yield the lowest total cost is _______ Ex. In whole number, use Q, <, <=, >, >= to specify the range)

c) The range of output for location alternative 3 that would yield the lowest total cost is ______ (Ex. In whole number, use Q, <, <=, >, >= to specify the range)

d) The alternative would yield the lowest total cost for an expected annual volume of 4005 boats is location _____ (Ex. type in as L1, L2, or L3)
Answer #1

40000 35000 30000 25000 Total cost 20000 TC1 = 7958 + 3.10 -TC2 = 5879 + 4.50 15000 TC3 = 11897 + 2.20 10000 5000 1485 4376.7

(a)

The range of output for location alternative 1 that would yield the lowest total cost is:

1485 <= Q <= 4376

(b)

The range of output for location alternative 2 that would yield the lowest total cost is:

0 <= Q <= 1485

(c)

The range of output for location alternative 3 that would yield the lowest total cost is

Q >= 4377

(d)

Q=4005 is in the range 1485 <= Q <= 4376. So, location alternative 1 will give the lest cost.